# Underwater Variable Zoom: Depth-Guided Perception Network for Underwater Image Enhancement

### v1

@@ III-B Dual-Attention Module
In complex underwater imaging processes, the depth-related degradation attenuation is quite common. As shown in Fig. 4, we design a DAM to generate reasonable depth maps, consisting of channel attention (CA) and self-attention (SA). In CA, we combine the convolutional results of two pooling operations to extract channel importance $f_{ca}$ from the encoding feature $f$:
$f_{ca}=f\times\left(AvgP(f)_{clc}+MaxP(f)_{clc}\right)_{sig},\$ (5)
where $AvgP(\cdot)$ and $AvgP(\cdot)$ represent the average pooling layer and maximum pooling layer, respectively, and $(\cdot)_{sig}$ is the sigmoid function. Next, we input the features annotated with channel weights into SA, allowing regions with different degradation degrees to have distinct weights. SA first uses the average pooling layer to reduce the scale, and then obtains spatial importance $f_{sa}$ by the following process:
$q=k=v=\left(f_{ca}\right)_{c},$
$f_{sa}=f_{ca}+\gamma\times\left((\mathrm{q}\times k)_{s}\times v\right),$ (6)
where $(\cdot)_{s}$ is a softmax function, and $\gamma$ is a learnable parameter used to adjust the intensity. We use the deconvolution $(\cdot)_{dc}$ to reshape features into input scale and annotate spatial weights back to input features. The hidden feature $f_{h}$ with regional degradation prominence is obtained by the fusion of two attention features:
$f_{h}=f+f_{ca}\times\left(f_{sa}\right)_{dc}$ (7)
@@
III-C Residual Block and Residual Skip Block
Figure 5: Structure of the Residual Block (RB) and Residual Skip Block (RSB)
Fig. 5 shows the details of RB and RSB. In both blocks, we extract features through convolution-activation-convolution structures and residual connection, while extra 1 × 1 convolution blocks are applied to RSB to fuse decoded features from DEN. We add many RB to solve the gradient vanishing, and assist the ASN in generating a more complete regression image $\widehat{X}$ by RSB.
III-D Training Strategy
@@
$L_{s}=\sqrt{||x-\left.y\right|^{2}+\varepsilon^{2}}+\lambda_{2}(1-\operatorname{SSIM}(x,y)),$ (9)
@@
IV Experiment
IV-A Implementation Details
@@
TABLE I: Metrics comparison of all methods on the CYCLE dataset (optimal, sub-optimal)
@@
IV-B Comparison on reference dataset
@@
Table I shows the quantitative comparison on the CYCLE dataset, where UVZ outperforms other state-of-the-art methods in all quality assessments, particularly with 11.9% and 8.9% performance improvements in PSNR and UIQM. Compared to other methods, Water-Net is more effective in removing color deviation, while SCNet notably enhances the red channel. Therefore, these methods obtain most sub-optimal metrics in Table I. However, in terms of runtime, our method does not exhibit outstanding speed. To accurately guide region perception, UVZ adopts a two-stage design and multiple attentions. While these configurations require substantial computing resources, the model performance is nonetheless impressive.
IV-C Comparison on no-reference dataset
We present enhancement examples from LNRUD, LSUI, UFO and UIEB datasets in Figs. 9-12. In Fig. 9 (a) and Fig. 10 (a), the color enhancement from the UVZ is the most pronounced and well-established. While other methods mitigate the effects of color deviation, blue-green color still dominates the results. We enlarged the fishes in the near-far scenes in Fig. 9 (b) and Fig. 10 (b), respectively. It can see that only TEBCF and UVZ effectively remove the blue-green color, but the dull color of the former reduces the attractiveness.
From Fig. 11 and Fig. 12, it can be observed that L2UWE exhibits bright contrast enhancement, but over-exposure affects the color. The enhanced results of TEBCF, WWPE and Water-Net have low brightness and poor clarity. CBM effectively enhances image details, while TACL, USUIR and PUGAN alleviate the impact of color deviation. However, blue-green colors still dominate in their enhanced images. Despite the significant saturation increase in SCNet and Spectroformer, the enlarged far-view Fig. 12 (c) still shows deviated colors. Benefiting from its depth perception ability, our UVZ has excellent performance in terms of color, saturation, and visibility, whether in near or far scenes.
TABLE II: Metrics comparison of all methods on four benchmark datasets (optimal, sub-optimal)
@@ IV-C Comparison on no-reference dataset
Table II gives the quantitative metrics on the above datasets, showcasing that our method excels in most metrics, especially the UICM representing the color metric. CBM exhibits impressive NIQE metrics, signifying a natural restoration of scene details. However, its color performance heavily affected the UICM evaluation.
IV-D Ablation Study
@@
TABLE III: Metrics comparison of the first stage configurations on the CYCLE dataset (optimal, sub-optimal)
@@
We conducted a detailed ablation study for two stages separately. First, we removed different components in the first stage, and fixed the second stage configuration to train the entire network. The ablation configurations are as follows: (1) -w/o DAM, representing the removal of DAM in DEN, retaining only skip connections. (2) -w/o RSB, representing the removal of feature transmission from RSB in ASN. (3) -w/o ASN, representing the removal of auxiliary supervision from ASN in the training process. Fig. 13 compares the effects of the above configurations. In the absence of DAM, the depth estimation of scene details lacks accuracy, and the removal of RSB introduces confusion in the predicted depth. Therefore, the enhanced images of two configurations suffer from excessive red channel enhancement. The depth image of -w/o ASN deviates from the underwater image, leading to inaccurate color enhancement. In Table III, although the referenced metrics of the complete network are slightly lower than that of -w/o ASN, the advantage of the no-reference metrics still proves a significant gain of ASN.
@@
TABLE IV: Metrics comparison of the second stage configurations on the CYCLE dataset (optimal, sub-optimal)
@@
Second, we fixed the configuration of the first stage, and removed different components in the second stage to obtain the following ablation network: (1) -w/o RB, representing the removal of all RB. (2) -w/o D, representing the removal of the depth maps in DPM, retaining only the second stage. (3) -w/o Rev, representing the removal of depth map reversal. (4) -w/o RN, representing the removal of RN in the R3N transformation, directing the local branch only through the reshaped depth map. (5) -w/o DPM, representing the removal of the DPM, retaining only skip connections. In Fig. 14. (a), the color of the complete network is the most consistent with GT. When the network lacks RB or DPM, the enhanced images are still affected by color deviation. While the red channel enhancement of - w/o D, -w/o Rev and - w/o RN is insufficient, and the color distribution of -w/o RN is chaotic. In the enlarged far water region of Fig. 14. (b), only the -w/o RN and the complete network effectively counter the color deviation.
IV-E Extended Applications
V Conclusion
This paper proposes a novel UIE framework for depth-guided perception. In DEN, we use DAM to focus on regions with severe degradation, making the predicted depth map more reasonable. To ensure the relevance of depth map to scene, ASN is introduced into the training process to generate regression image. In addition, we design a DPM to capture near-far features, with the R3N transformation guiding the fusion of non-local and local branches. Extensive experiments on benchmark datasets demonstrate the advantages of our method and the gains of each module. Moreover, our method proves effectiveness across different visual tasks, improving visibility in challenging lighting conditions. In future work, we aim to optimize model efficiency and explore scenario enhancement by multiple semantic information.

### v3

@@ 3.3 Dual-Attention Module
In complex underwater imaging processes, the depth-related degradation attenuation is quite common. As shown in Fig. 5, we design a DAM that includes channel and spatial monitoring mechanisms to generate reasonable depth maps. First, DAM compresses the encoded feature $f$ into two channel descriptors through different pooling operations. Then, the channel dependence used to adapt the feature representation is captured by the squeeze-excitation process [40]. The above process can be represented as follows:
$f_{c}=f\times\left(AvgP(f)_{clc}+MaxP(f)_{clc}\right)_{sig},\$ (5)
where $AvgP(\cdot)$ and $AvgP(\cdot)$ represent the average pooling layer and maximum pooling layer, respectively, and $(\cdot)_{sig}$ is the sigmoid function. The features annotated with channel weights $f_{c}$ is input into the next stage, allowing regions with different degradation degrees to have distinct weights. First, we use the average pooling layer to reduce the feature scale, and then capture the spatial dependence by the query vector $q$, the key vector $k$, and the value vector $v$. In this process, we introduce a learnable parameter $\gamma$ to dynamically adjust the learning intensity:
$q=k=v=\left(AvgP(f_{c})\right)_{c},$
$f_{s}=f_{c}+\gamma\times\left((\mathrm{q}\times k)_{soft}\times v\right),$ (6)
where $(\cdot)_{soft}$ is a softmax function, and $\gamma$ is a learnable parameter used to adjust the intensity. We use the deconvolution $(\cdot)_{dc}$ to reshape features into input scale and annotate spatial weights $f_{s}$ back to input features. We fused the two attention features and perform residual connections with the original input $f$. As a result, the proposed model can capture the hidden feature $f_{h}$ with regional degradation prominence:
$f_{h}=f+f_{c}\times\left(f_{s}\right)_{dc}$ (7)
@@
3.4 Training Strategy
@@
$L_{s}=\sqrt{||x-\left.y\right||^{2}+\varepsilon^{2}}+\lambda_{2}(1-\operatorname{SSIM}(x,y)),$ (9)
@@
4 Experiment
4.1 Implementation Details
@@ 4.2 Comparison on reference dataset
Table 1: Metrics comparison of all methods on the CYCLE dataset (optimal, sub-optimal)
@@
Table 1 shows the quantitative comparison on the CYCLE dataset, where UVZ outperforms other state-of-the-art methods in all quality assessments, particularly with 11.9% and 8.9% performance improvements in PSNR and UIQM. Compared to other methods, Water-Net is more effective in removing color deviation, while SCNet notably enhances the red channel. Therefore, these methods obtain most sub-optimal metrics in Table 1. However, in terms of runtime, our method does not exhibit outstanding speed. To accurately guide region perception, UVZ adopts a two-stage design and multiple attentions. While these configurations require substantial computing resources, the model performance is nonetheless impressive.
4.3 Comparison on no-reference dataset
We present enhancement examples from LNRUD, LSUI, UFO and UIEB datasets in Figs. 9-12. In Fig. 9 and Fig. 10, the color enhancement from the UVZ is the most pronounced and well-established. While other methods mitigate the effects of color deviation, blue-green color still dominates the results. We enlarged the fishes in the near-far scenes in Fig. 9 (c) and Fig. 10 (c), respectively. It can see that only TEBCF and UVZ effectively remove the blue-green color, but the dull color of the former reduces the attractiveness.
From Fig. 11 and Fig. 12, it can be observed that L2UWE exhibits bright contrast enhancement, but over-exposure affects the color. The enhanced results of TEBCF, WWPE and Water-Net have low brightness and poor clarity. CBM effectively enhances image details, while TACL, USUIR and PUGAN alleviate the impact of color deviation. However, blue-green colors still dominate in their enhanced images. Despite the significant saturation increase in SCNet and Spectroformer, they still show deviated colors in the far scenes. Benefiting from its depth perception ability, our UVZ has excellent performance in terms of color, saturation, and visibility, whether in near or far scenes.
@@ 4.3 Comparison on no-reference dataset
Table 2: Metrics comparison of all methods on four benchmark datasets (optimal, sub-optimal)
@@
Table 2 gives the quantitative metrics on the above datasets, showcasing that our method excels in most metrics, especially the UICM representing the color metric. CBM exhibits impressive NIQE metrics, signifying a natural restoration of scene details. However, its color performance heavily affected the UICM evaluation.
4.4 Ablation Study
We conducted a detailed ablation study for two stages separately. First, we removed different components in the first stage, and fixed the second stage configuration to train the entire network. The ablation configurations are as follows: (1) -w/o DAM, representing the removal of DAM in DEN, retaining only skip connections. (2) -w/o RSB, representing the removal of feature transmission from RSB in ASN. (3) -w/o ASN, representing the removal of auxiliary supervision from ASN in the training process. Fig. 13 compares the effects of the above configurations. In the absence of DAM, the depth estimation of scene details lacks accuracy, and the removal of RSB introduces confusion in the predicted depth. Therefore, the enhanced images of two configurations suffer from excessive red channel enhancement. The depth image of -w/o ASN deviates from the underwater image, leading to inaccurate color enhancement. In Table 3, although the referenced metrics of the complete network are slightly lower than that of -w/o ASN, the advantage of the no-reference metrics still proves a significant gain of ASN.
@@ 4.4 Ablation Study
Table 3: Metrics comparison of the first stage configurations on the CYCLE dataset (optimal, sub-optimal)
@@
Second, we fixed the configuration of the first stage, and removed different components in the second stage to obtain the following ablation network: (1) -w/o RB, representing the removal of all RB. (2) -w/o D, representing the removal of the depth maps in DPM, retaining only the second stage. (3) -w/o Rev, representing the removal of depth map reversal. (4) -w/o RS, representing the removal of RS in the R3S transformation, directing the local branch only through the reshaped depth map. (5) -w/o DPM, representing the removal of the DPM, retaining only skip connections. In Fig. 14. (a), the color of the complete network is the most consistent with GT. When the network lacks RB or DPM, the enhanced image is still affected by color deviation. While the red channel enhancement of - w/o D, -w/o Rev and - w/o RS is insufficient, and the color distribution of -w/o RS is chaotic. In the enlarged far water region of Fig. 14. (b), only the -w/o RS and the complete network effectively counter the color deviation. In Table 4, the complete network is weaker than the -w/o RS in UIQM, but it outperforms the other configured networks in the remaining four metrics.
@@
Table 4: Metrics comparison of the second stage configurations on the CYCLE dataset (optimal, sub-optimal)
@@ 4.5 Extended Applications
5 Conclusion
This paper proposes a novel UIE framework for depth-guided perception. In DEN, we use DAM to focus on regions with severe degradation, making the predicted depth map more reasonable. To ensure the relevance of depth map to scene, ASN is introduced into the training process to generate regression image. In addition, we design a DPM to capture near-far features, with the R3S transformation guiding the fusion of non-local and local branches. Extensive experiments on benchmark datasets demonstrate the advantages of our method and the gains of each module. Moreover, our method proves effectiveness across different visual tasks, improving visibility in challenging lighting conditions. However, UVZ exhibits limited region division in response to diverse underwater scenes, and lags behind lightweight models in speed. In future work, we will aim to optimize the model efficiency and explore multiple semantic information fusion to enhance scenario understanding.